### v4

# Re(l)ality:

Abstract
Using the fibre bundle framework, this work investigates the conceptual and mathematical foundations of reference frames in General Relativity by contrasting two paradigms. The View from Nowhere interprets frame representations as perspectives on an invariant equivalence class, while the View from Everywhere posits each frame representation as constituting a fully-fledged reality itself. What emerges is a conception of reality that I term ”Relality.” The paper critically examines the philosophical and practical implications of these views, with a focus on reconciling theory with experimental practice. Central to the discussion is the challenge of providing a perspicuous characterisation of ontology. The View from Nowhere aligns with the so-called ‘ sophisticated approach to symmetries’ and complicates the empirical grounding of theoretical constructs. In contrast, the View from Everywhere offers a relational ontology that avoids the abstraction of equivalence classes.
1 Introduction
In theoretical frameworks known as gauge theories, such as General Relativity (GR), the concept of observables refers to quantities that remain unchanged under specific transformations called gauge transformations (Dirac, 1964). These are transformations that do not alter the physical content of the theory, also referred to as the ‘observable content’. Consequently, gauge transformations are often understood as mere mathematical re-expressions of the theory’s formalism and are seen as indicators of a representational redundancy in the theory, where multiple mathematical descriptions correspond to the same physical reality.111‘The ‘gauge argument’ is well known in both physical and philosophical literature (see Berghofer et al. (2023) for a recent overview). I will therefore say no more on the subject, assuming the reader is familiar with the main concepts summarised above. As a final point, it should be pointed out that the presence of gauge symmetries is closely related to the threat of indeterminism. It is precisely to save the theory from this threat that the concept of gauge was introduced in Dirac (1964).
The standard assumption in physics is that a procedure of ‘gauge-fixing’ is necessary to remove the descriptive redundancy, ensuring a unique description of physical configurations through the definition of observable quantities. [^2] It is interesting that the verb ‘fix’ in English can mean two things, both of which are correct in this context: (i) adjusting the formalism to eliminate redundant degrees of freedom; (ii) setting a specific condition (called the gauge condition) from a shortlist of possible conditions, in order to construct a single observable quantity from redundant degrees of freedom. Along these lines, this paper explores how observables are constructed and interpreted in GR.
For a theory to achieve empirical validity, its gauge-invariant observables must correspond to measurable quantities, facilitating comparisons between theoretical predictions and experimental observations. This correspondence is a critical criterion for the acceptance of any scientific theory. [^3] Although a theory cannot be entirely confirmed or falsified by experimental evidence, its ability to allow meaningful comparisons with reality is essential to avoid being labeled as ad hoc (Leplin, 1975).
All measurements are inherently local, as they rely on observations within specific regions of spacetime. Much of our experience involves measuring locally defined variables at particular spacetime points or within localised regions. As Gary and Giddings (2007) emphasise, “all we can truly observe is localised — we have no access to infinity.” In GR, gauge symmetries are expressed as diffeomorphisms. As a result, any quantity defined locally in spacetime does not remain invariant under these transformations, raising challenges for constructing local, gauge-invariant observables.
In a diffeomorphism-invariant theory, variables can be broadly categorized into two types: (1) highly non-local quantities, which are defined across the entire spacetime, such as integrals over global regions, and (2) relational quantities, which are constructed by correlating field values. This work focusses primarily on the second category, commonly referred to as relational observables.
A recent approach to constructing local observables in GR was introduced by Rovelli (2002b) and formalised by Dittrich (2006, 2007). This approach involves correlating two partial observables—quantities that are individually gauge-variant—to form a gauge-invariant complete observable. These complete observables embody a relational notion of locality: instead of being situated within a fixed spacetime background, physical states are defined in relation to other fields. In simpler terms, a complete observable represents the value of one partial observable when the value of another is specified.
This relational framework is closely connected to the role of reference frames in GR, as discussed in Bamonti (2023). Reference frames offer a structured way to define gauge-invariant quantities relationally. In practice, constructing relational observables often involves treating one partial observable as a spatiotemporal reference frame for another, thereby highlighting the inherently relational nature of locality in this framework.
As Bamonti (2023); Bamonti and Gomes (2024b) argue, one way to construct local, complete observables in GR is through dynamically coupled reference frames. Specifically, when a pair of partial observables forms a valid solution to the dynamics of the theory, only the diagonal action of a dynamical symmetry (such as a diffeomorphism in GR) preserves this solution.
This paper addresses a central question: how should we interpret two local observables expressed in terms of distinct reference frames? Two primary views aim to answer this question.
The first, the View from Nowhere is the dominant perspective in the literature. It often asserts that changing a reference frame does not affect the physical content of the description, as a reference frame provides merely a perspective on a shared, objective reality. This idea has its roots in Special Relativity, where different inertial reference frames offer distinct perspectives on a Lorentz-invariant physical reality. Prominent examples of this view include the perspective-neutral framework introduced by Vanrietvelde et al. (2020) and related work by Giacomini et al. (2019) and Kabel et al. (2024). Explicitly, Kabel et al. (2024) state “if we assume that the covariance of physical laws under classical changes of reference frame extends linearly to encompass changes of QRF, then the latter should not affect the physical situation. Thus, whether a system is in superposition or is entangled, and how these properties change dynamically, becomes a mere matter of perspective” [my italics].
Importantly, throughout the paper I will use the term ‘perspective’ in the sense also adopted in (Adlam, 2024b) and there called ‘physical perspectivalism’: a perspective is defined by physical observers which play the role of a reference frame. [^4] The term ’perspective’ is widely used in the philosophy of science. For example, Michela Massimi uses the notion of perspective in terms of a historically and culturally situated epistemic lens within a debate on scientific realism Massimi (2022). Another possibility is to use the term perspective only in an epistemic way: a perspective merely indicates the degree of partial knowledge of a conscious observer. This form of perspectivalism is known as ‘epistemic perspectivalism’.
It is worth clarifying the connection I am drawing between the View from Nowhere and Perspectivalism. In the literature, e.g. in Adlam (2024b), it is argued that perspectivalism rejects the View from Nowhere and that “perspectivalists are fond of the mantra ‘there is no view from nowhere’ ” (ivi, p.4). [^5] See also Ruyant (2020): “(various different forms of perspectivalism) share the general idea that there is no ‘view from nowhere’ ”. However, I believe this claim stems from a misuse of the term ‘perspective’. In particular, I argue that a perspective is always a perspective on something which stays unchanged. The perspective neutral structure presented in Vanrietvelde et al. (2020) is exactly this ‘something’ on which various perspectives are defined.
In characterising physical perspectivalism, Adlam advocates a moderate version, whereby perspective-neutral physical facts exist, but have no empirical meaning. Well, moderate perspectivalism advocates the existence of a View from Nowhere, only it does not give it empirical meaning, but it is there.666‘There’ where? I claim nowhere! See fn.22 in §3. As Adlam herself also claims:
It appears that considerations from quantum mechanics and relativity do provide support for the idea that empirically meaningful descriptions typically have to be relativized to a perspective. But this in and of itself does not tell us whether we should adopt mod- erate or strong physical perspectivalism. To make that choice, we must decide whether these considerations suggest that all facts about physical reality must be relativized to a perspective, or whether they instead support the existence of some perspective-neutral facts about physical reality, over and above descriptions relativized to perspectives (Adlam, 2024b, p.6).
In short, moderate perspectivalism understands each perspective as a partial, limited glimpse into a more complete reality, as is also evident from Adlam’s admission that moderate perspectivalism is motivated by the epistemic humility sustained by epistemic perspectivalists. A moderate perspectivalist can thus be seen as the proponent of a vision of a kind of ‘ontological humility’.
The strong version of perspectivalism, on the other hand, rejects the existence of any perspective-neutral fact: there are only perspectives. Based on what I argued earlier, this version of perspectivalism is a non-sense by definition: a perspective is always a perspective on something. Consequently, the perspectivalism that will be criticised in this paper in association with the View from Nowhere is the moderate version: the only one that makes sense.777 As will become clear later in the introduction, another critical point of Adlam’s cited work is that she does not distinguish between independence and freedom, but only adopt the quite ambiguous term ‘neutral’ and this leads her to consider that the presence of connections between perspectives and strong perspectivalism cannot be compatible. However, if the connections are considered perspective-independent facts, there is no incompatibility.
The View from Everywhere, proposed in this paper, challenges the assumption of an underlying objective reality. I argue that there is no direct analogy between GR and Special Relativity in this context. While Special Relativity relies on a fixed, Lorentz-invariant framework which proves us with a “gods-eye’ description of the way in which all of the perspectives are related” (van Fraassen, 2008), GR introduces local gauge-invariant quantities that are not necessarily frame-invariant. Local relational observables, although gauge-invariant, can differ numerically when expressed in different reference frames. Consequently, the View from Everywhere suggests that frame representations are not merely perspectives on an objective reality; they are fundamental realities themselves. No God’s view is allowed.
To clarify the terminology used in this work, I distinguish between spatiotemporally explicit and implicit quantities, as well as local and non-local quantities. Additionally, the term ”invariant” requires careful consideration, particularly in distinguishing between independence and freedom.
Table 1 provides an overview of these distinctions, presenting a structured summary of the various types of quantities in GR along with their defining characteristics. [^8] Notice that a metric $g_{\mu\nu}(x^{\rho})$ written in some coordinate system $\{x^{\rho}\}\in\mathbb{R}^{N}$ is a frame-free, spatiotemporally explicit, local, and gauge-variant object.
Table 1: Summary of distinctions between different types of quantities that can be defined in GR.
$g_{ab}(p)$ Spatiotemporally explicit and local Frame free Gauge-variant
[…]
$[g_{ab}]$ Spatiotemporally implicit Frame free Gauge-invariant
$g_{IJ}(\phi)$ Frame explicit and local Frame dependent Gauge-invariant or not, depending on $\phi$
[…]
$[g_{IJ}]$ Frame implicit Frame independent Gauge-invariant
Within the framework of non-relational quantities, it is possible to differentiate between quantities that are:
  1. i)
    Spatiotemporally explicit: A quantity is said to be spatiotemporally explicit if it is defined by a single function of a given region $U\subseteq\mathcal{M}$. They can be distinguished into:
    1. i.a)
      Spatiotemporally local: A spatiotemporally local quantity is one that is determined solely by the properties at a single point, understood as a region of $\mathcal{M}$ having a zero Lebesgue measure. It is “pointwise” defined and does not depend on information from neighboring points. For example, the metric tensor $g_{ab}(p)$ is defined locally on the manifold—it provides the geometric structure exactly at $p$. Such objects are not gauge-invariant.
    2. i.b)
      Spatiotemporally non-local: A spatiotemporally non-local quantity, on the other hand, is defined over an extended region of $\mathcal{M}$. Such quantities involve integrals or sums over areas or volumes, meaning that their values cannot be determined solely by the properties at an isolated point. For instance, the total volume obtained by integrating the metric over a region is non-local because it aggregates information from an extended domain of $\mathcal{M}$. Such objects are gauge-invariant. [^9] Strictly speaking, these quantities are gauge-invariant if the integration extremes are not fixed. Otherwise, also a volume varies even under a diffeomorphic transformation. See, e.g., formulas 15 and 17 in Rovelli (2014).
  2. ii)
    Spatiotemporally implicit: A quantity is spatiotemporally implicit if it cannot be represented by a single unique function on $\mathcal{M}$, but only as an equivalence class. Such quantity is usually denoted using square brackets: for example, $[g_{ab}]$. Such objects are gauge-invariant.
Introducing reference frames allows us to construct relational quantities. Here, I provisionally denote a reference frame without further specification as a set of four scalar quantities $\phi^{(I)}\mid_{I=1,\dots,4}$ uniquely associating a real quadruple $x\in\mathbb{R}^{4}$ with each point $p\in\mathcal{M}$ (see the next section for a more detailed discussion; see Bamonti (2023) for a comprehensive discussion). Similarly to the above, it is possible to distinguish different definitions of relational objects in GR, when written in terms of reference frames.
We can distinguish between relational quantities that are:
  1. I)
    Frame explicit: These quantities are described by a single function of a given region ‘covered’ by the reference frame field (we can safely imagine a reference frame exactly as ‘chart’ in differential geometry). They can be subdivided into:
    1. I.A)
      Frame local: A frame local quantity is one that is determined solely by the properties at a single point $p:=\phi^{-1}(x)$ – for example, $g_{IJ}(\phi)$ (see the next section for the detailed construction and meaning of notation). [^10] One might wonder whether $g_{IJ}(\phi)$ is not a functional, since each $\phi$ is a scalar field (a functions of point $p\in\mathcal{M}$). However, even though $\phi^{(I)}(p)$ are scalar functions defined on the manifold, when we write $g_{IJ}(\phi)$ we are expressing the metric components in terms of the values provided by each $\phi$ at each point. In other words, each point $p$ is mapped to a coordinate $x=\phi(p)\in\mathbb{R}^{4}$, and $g$ is a function that assigns the corresponding metric tensor $g_{IJ}$ to each such $x$. This is different from a functional, which would take an entire function as input and yield a number. Thus, $g_{IJ}(\phi)$ is a function, not a functional. Its value at a point $p$ depends on the local value of the reference fields at that point, so a frame-local quantity is defined by ‘pointwise’ information alone, where the notion of locality is relational in terms of other fields and not manifold points. These objects are gauge-invariant when the reference frame is appropriately coupled to the metric. However, if the reference frame is uncoupled, the quantity may remain relational without being fully gauge-invariant (see Bamonti and Gomes (2024b)).
    2. I.B)
      Frame non-local: A frame non-local quantity is defined over an extended region parametrised by the reference frame. For example, total volume integral of a given gauge-invariant metric $g_{IJ}(\phi)$ locally defined is a frame non-local quantity, since its value depends on the metric values over a range of points rather than at a single point. Also such objects are gauge-invariant only if the reference frame is coupled with the metric. Otherwise, we can independently apply a diffeomorphism either to the integration extremes or to the metric and the situation is analogous to that in which after a diffeomorphism the metric (the integrand) changes, but the extremes stay fixed, resulting in the integral value changing (see fn.9 above).
  2. II)
    Frame implicit: A frame implicit quantity is not expressible by a unique function, but only as a collection of frame-explicit quantities: for example an equivalence class. In such a case, the equivalence class can still be denoted using square brackets as $[g_{IJ}]$ and constitutes a gauge-invariant object. See §4 for a discussion on such object.
NB: of course, a frame explicit object will be also spatiotemporally explicit, in the specific sense enclosed by the relational locality inherent to the use of reference frames. To avoid confusion, I have decided to distinguish between frame-explicit and spatiotemporally explicit objects, preserving the term ‘spatiotemporal’ to indicate objects written in terms of manifold points (if it helps, the reader may also interpret a spatiotemporally explicit object as ‘manifold explicit’). [^11] The concept of spacetime lacks a universally agreed definition, and its interpretation varies across different frameworks. It may be understood as: (i) the manifold $\mathcal{M}$; (ii) the combination $(\mathcal{M},g_{ab})$, which consists of the manifold equipped with the metric field representing the gravitational field; or (iii) the gravitational field $g_{ab}$ alone. In interpretations (i) and (ii), the distinction lies in whether $\mathcal{M}$ is considered an ontologically independent entity—a stage upon which dynamical variables act—or whether it is inseparable from the fields. Interpretation (iii), however, treats $\mathcal{M}$ as a purely mathematical construct without ontological status (see Rovelli and Gaul (2000); Rovelli (2006); Rovelli and Vidotto (2015); Einstein et al. (2015)). An additional perspective considers spacetime as an emergent structure, defined in terms of observables in the Dirac (gauge-invariant) sense. These observables correspond to the happening of events, which specify the ”when” and ”where” of physical phenomena in a relational manner. Accordingly, the ‘where and when’ are consequential to the happening of an event (which is a gauge-invariant observable). So it is the happening of the event that determines where and when it happens and not the event happening at a where and when. In this paper I choose the more conservative option and identify space-time with option (ii).
Finally, following and expanding Wallace (2019)’s work, in which he distinguishes between the concepts of (coordinate) independence and (coordinate) freedom, I distinguish between:
  1. 1.
    Reference frame-dependence: a quantity is reference frame-dependent if its definition depend on the reference frame in which is defined.
  2. 2.
    Reference frame-independence: a quantity is reference frame-independent if its definition does not depend on the reference frame in which it is, however, defined.
  3. 3.
    Reference frame-freedom: a quantity is reference frame-free if does not need any reference frame to be defined.
The following sections delve deeper into these ideas, exploring the implications of reference frame dependence and offering a comprehensive framework to interpret relational observables in GR. Furthermore, the introduction of the View from Everywhere allows me to provide a perspicuous characterisation of relational ontology (section 4), which is not readily provided in the View from Nowhere (section 3).
2 The Fibre Bundle Formalism: A Gauge Perspective on Reference Frames
In this section, I use the fibre-bundle formalism to describe reference frames and relational observables. This approach, widely used in foundational studies of gauge theories (see e.g. Healey (2007); Weatherall (2016)), is primarily attributed to the work of Gomes (see, e.g., Gomes (2023a, b) for a rigorous treatment). Let $M$ represent the space of models $m$ of the theory. The space $M$ can be described as a principal bundle with $S$ as its structure group and $[M]:=\{[m]\mid m\in M\}$ as its base space, where $[m]$ identifies the equivalence class of models under the transformations of $\mathcal{S}$. In this formalism, selecting a reference frame involves defining a unique section map $\sigma:[m]\to\sigma([m])\in M$, which smoothly maps equivalence classes of models to individual models in the space . This corresponds to choosing a submanifold in the fibre bundle that intersects each fibre $\mathbb{F}_{m}:=\mathrm{pr}^{-1}([m])$ exactly once, with $\mathrm{pr}:m\to[m]$ being the projection map.
Choosing a reference frame can also be interpreted as selecting a specific gauge. In fact, the choice of a gauge is equivalent to the choice of a reference frame (see Dittrich (2007) for insights on the relationship between gauge-fixed observables and relational observables. See also Bamonti and Thébault (2024) for an example of such relationship in FLRW cosmology). In gauge theories with a principal fibre bundle structure, gauge-fixing determines a section through the fibre bundle. Given a symmetry group $S$, each fibre corresponds to a gauge orbit — the set of all configurations of a field that are related by gauge transformations, generated by the constraints of the theory. Specifically, there exists a one-to-one correspondence between each gauge orbit and the equivalence class of models under the symmetry $s\in S$. More concretely, for $m\in M$ and $s\in S$, the orbit $\mathcal{O}_{m}$ is defined as $\mathcal{O}_{m}=\{m_{s}\mid s\in S\}$, where $m_{s}$ denotes the model to which the transformation $s$ is applied. Due to the symmetry of the theory, the model space includes significant redundancy. The physical content of the theory is captured by a single representative from each gauge orbit, while other models within the same orbit are redundant, isomorphic copies. [^12] This does not imply that gauge freedom constitutes mere ”descriptive fluff” (see Earman (2004)). On the contrary, the correspondence between gauge-fixing and reference frame selection highlights the relational nature of physics. The additional degrees of freedom are meaningful as they represent the possible ways a system can form observables relative to another system. Isomorphic models provide ”handles” through which systems can couple; see Rovelli (2014) and Adlam (2024a). This redundancy is typically resolved by fixing a gauge. Similarly, in GR, one fixes a reference frame through a coordinate gauge condition (see Bamonti (2023); Gomes (2024b)).
In the following, I provide a concrete example to illustrate the reference frame formalism and the gauge-fixing procedure within the bundle framework. This example serves as a foundation for discussing the two distinct perspectives on interpreting local observables in GR.
Let the space of models be $M=\text{Lor}(\mathcal{M})$. This denotes considering tuples $\langle\mathcal{M},g_{ab}\rangle$ as possible models, focusing on ‘vacuum’ GR. [^13] A typical model $M\ni m=\langle\mathcal{M},g_{ab},\phi\rangle$ consists of a manifold $\mathcal{M}$, a (Lorentzian) metric $g_{ab}$, and some matter field $\phi$. The set $\text{Lor}(\mathcal{M})$ can be interpreted as a principal bundle with $\mathcal{S}=\text{Diff}(\mathcal{M})$ as its structure group and $[\text{Lor}(\mathcal{M})]:=\{[g_{ab}],g_{ab}\in\text{Lor}(\mathcal{M})\}$ as its base space. [^14] Field theories like GR face challenges in defining a $[\text{Lor}(\mathcal{M})]\times\text{Diff}(\mathcal{M})$ product structure, even locally ( this is consistent with the general impossibility of defining a global reference frame). This is viable only for globally hyperbolic spacetimes admitting a CMC foliation. Furthermore, the Gribov obstruction limits the construction to a local product structure (Gribov (1978); Henneaux and Teitelboim (1994). Each equivalence class $[g_{ab}]$ consists of diff-related metrics:
$[g_{ab}]:=\{g_{ab},(d^{*}g)_{ab},\dots\}.$
Suppose we construct four scalar quantities, $\Re^{(I)}_{g}$, $I=1,\ldots,4$, from the metric $g$. These are known as Kretschmann-Komar scalars, named after Kretschmann (1918) and Komar (1958) (see also Bergmann and Komar (1960, 1962)). [^15] Komar (1958)) derived these scalars using an eigenvalue problem involving the Riemann tensor $R_{abcd}$ and an anti-symmetric tensor $V_{cd}$: $R_{abcd}-(g_{ac}g_{bd}+g_{ad}g_{bc})V_{cd}=0$. The set $\{\Re^{(I)}_{g}\}$ provides a spatiotemporal reference frame for the metric field itself, aligning with the relational strategy: “Rather than fixing an observable at specific coordinates, its location is defined relative to features of the state” (Harlow and qiang Wu, 2021).
A reference frame can be defined as a physical system yielding a local diffeomorphism:
$\Re_{g}^{(I)}:=(\Re_{g}^{(1)},\cdots\Re_{g}^{(4)}):U\subseteq\mathcal{M}\rightarrow\mathbb{R}^{4},$ (1)
which uniquely assigns four numbers to each point in $U$. Using this frame, tensors like the metric $g_{ab}$ can locally be ‘coordinatised’ by $\{\Re^{I}_{g}\}$. Specifically, for all isomorphic models, the gauge-invariant relational observable
$g_{IJ}(\Re_{g}):=\left[\Re_{g}^{-1}\right]^{*}g_{ab}$
produces a set of 10 scalar functions indexed by $I$ and $J$, constructed from the metric tensor and its derivatives. [^16] Viable reference frames $\Re_{g}$ must be locally invertible. In spacetimes with continuous symmetries, such as metrics admitting Killing vectors, this condition may fail, making the scalars linearly dependent. Thus, linear independence of $\{\Re_{g}^{(I)}\}$ is necessary for their viability as a reference frame. We should distinguish linear (in)dependence from functional (in)dependence. For example, one could have zero physical degrees of freedom — that is functional dependence — and still have a viable reference frame. Its gauge-invariance follows from the chain rule for the transformation of $\Re_{g}$:
$\forall d\in Diff(\mathcal{M}),\quad[\Re_{g}^{-1}]^{*}g_{ab}=[\big{(}d^{*}\Re_{g}\big{)}^{-1}]^{*}(d^{*}g)_{ab}.$ (2)
Thus, using this quadruple, we achieve a unique, gauge-invariant metric representation. Given initial data for $g_{ab}$, the dynamical evolution of $g_{IJ}(\Re_{g})$ is uniquely determined because $\Re_{g}^{(I)}$ are dynamically coupled to $g_{ab}$ — being functions of the metric itself. Consequently, diffeomorphisms act diagonally to preserve solutionhood: if $(g_{ab},\Re_{g}^{(I)})$ is a possible solution, then only $\big{(}(d^{*}g)_{ab},d^{*}\Re_{g}^{(I)}\big{)}$ is a still a possible solution.
This encapsulates the concept of a ’relational, gauge-invariant observable’ and a ’reference frame’. [^17] Strictly, a ‘relational observable’ is not automatically gauge-invariant (Bamonti and Gomes, 2024b).
The choice of $\Re_{g}$ as a reference frame can be formalised through a choice of a gauge. This allows us to fix a reference frame by a condition — valid for all the isomorphic models — that the models satisfy. This is most directly accomplished by postulating some constraint $F_{\Re_{g}}\in C^{\infty}(\mathcal{M})$, such that:
$\forall g_{ab}\in Lor,\exists!f_{\Re_{g}}\in Diff(M)\quad|\quad F_{\Re_{g}}(g_{ab})=0,$ (3)
where $f_{\Re_{g}}$ acts as the equivalent of a section map within the fibre-bundle framework. Specifically, this diffeomorphism $f_{\Re_{g}}:g_{ab}\to f_{\Re_{g}}^{*}g_{ab}$ serves as a projection operator, uniquely mapping any element of a given fibre to the section’s image. It is also called the projection operator for the section. [^18] The projection corresponding to the choice of a reference frame can be written down in terms of coordinate charts. For instance, in GR two common gauge-fixings are the De Donder gauge in the Lagrangian sector which corresponds to the condition $F(g)=\partial_{\mu}(g^{\mu\nu}\sqrt{g})=0$, and the CMC gauge in the Hamiltonian sector: $F(h^{ij},\pi_{ij})=h^{ij}\pi_{ij}=const$. The former corresponds to the use of coordinates that satisfy a relativistic wave equation $\Box x^{\mu}=0$. The latter, selects global simultaneity homogeneous 3-hypersurfaces $\Sigma_{\tau}$, parametrised by a universal time $\tau$. Notice that both are only partial gauge-fixings. More precisely, $f_{\Re_{g}}$ is the embedding map, acting within a fibre, from the fibre bundle manifold of models $Lor(\mathcal{M})$ to the image of the section map, and it is characterised by the auxiliary condition $F_{\Re_{g}}(g_{ab})=0$. The constraint $F_{\Re_{g}}(g_{ab})=0$ defines a ‘level surface’ of the section map along the fibres, effectively making the choice of a reference frame (or section) analogous to a gauge-fixing procedure. See Figure 1.191919 According to the condition (3), the choice of a reference frame, seen as the choice of a section, is based on the imposition of a set of conditions that a models must satisfy. This procedure is analogous to what Gauss (1902) proposed in order to describe a surface embedded in an ambient space not from an external point of view, i.e. using the coordinates of the ambient space, but ‘standing on the surface itself’. Such an embedded surface is intrinsically describable using some parametric equations. For example, a generic $n$-dimensional surface $\Sigma\subset N$, embedded in a generic $(n+1)$-dimensional Euclidean ambient space $N$ (characterised by $n+1$ coordinates $x^{i}$), can be described by $n$ parameters $u^{\alpha}$ with parametric equations of the type: $x^{i}(u^{\alpha})=0$. In our case, such equations take the form $F_{\mathcal{R}_{g}}(g_{ab})=0$, with $F_{\mathcal{R}_{g}}\in C^{\infty}(\mathcal{M})$ being a smooth and regular function.  [^20] Given a general symmetry group, its gauge orbits are in general not one-dimensional. The representation of figure 1 is faithful only for one-dimensional groups, whose action can be depicted in a one-to-one manner along the one-dimensional orbits.
Figure 1: The space of models $Lor(\mathcal{M})$ with its gauge group $Diff(\mathcal{M})$. Each point corresponds to a particular metric $g_{ab}$. A reference frame $\Re_{g}$ picks out a unique representative $(g_{ab})_{F_{\Re_{g}}}$ for each fibre $\mathbb{F}_{g}$. This is achieved via the projection map $f_{\Re_{g}}$ which projects a model within a fibre on the intersection between the fibre and a choice of section, whose ‘level surface’ is represented by $F_{\Re_{g}}$. Models belonging to the same fibre are taken to be physically equivalent, since a fibre corresponds to a gauge orbit. The space of equivalence classes of metric $[Lor(\mathcal{M})]:=\{[g_{ab}],g_{ab}\in Lor(\mathcal{M})\}$ is also referred to as the physical state space.
Given any doublet $\big{(}g_{ab},\Re_{g}^{(I)}\big{)}$, the action of $f_{\Re_{g}}$ will take that doublet to the unique and gauge-invariant reference frame representation of the metric $\big{[}\Re_{g}^{-1}\big{]}^{*}g_{ab}$.
This setup allows straightforward recovery of the analogue of equation (2) that demonstrate the gauge invariance of $f_{\Re_{g}}^{*}g_{ab}$
Proof.
From the diagonal action of the diffeomorphisms $d$, we have:
$f^{*}_{\Re_{g}}g_{ab}=f^{*}_{d^{*}\Re_{g}}\big{(}(d^{*}g)_{ab}\big{)},\forall d\in Diff(\mathcal{M}).$ (4)
Now, let me define
$(g_{ab})_{F_{\Re_{g}}}:=f^{*}_{\Re_{g}}g_{ab}\equiv\big{[}\Re_{g}^{-1}\big{]}^{*}g_{ab}\equiv g_{IJ}(\Re_{g}).$
The previous two equations imply that:
$\big{(}(d^{*}g)_{ab}\big{)}_{F_{d^{*}\Re_{g}}}=f^{*}_{d^{*}\Re_{g}}\big{(}(d^{*}g)_{ab}\big{)}=(g_{ab})_{F_{\Re_{g}}}.$
[…]
The map $f_{\mathcal{R}_{g}}$ is clearly model-dependent, but for each equivalence class of models $[g_{ab}]$ every model in this equivalence class will give rise to the same relational observable $f^{*}_{\mathcal{R}_{g}}g_{ab}$, since $f^{*}_{\mathcal{R}_{g}}g_{ab}=f^{*}_{\mathcal{R}_{d^{*}g}}[d^{*}g]_{ab}$, as shown in equation (4).
This framework explicitly ensures gauge invariance, although it remains dependent on the choice of section or reference frame. Consequently, it is not frame-invariant.
As I will show in the next section, this is why a choice of a reference frame is commonly labelled as a perspective on an equivalence class, within what I call the View from Nowhere.
3 The View from Nowhere: Frame representations are perspectives on an equivalence class
Naturally, there is nothing inherently special about $\{\Re^{(I)}_{g}\}$; any reference frame that provides a specific mapping $f$ for each isomorphism class suffices. The Kretschmann-Komar scalars are significant because of their explicit dynamical coupling to the metric. Crucially, the gauge invariance of the observable $g_{IJ}(\Re_{g})$ depends on this dynamical coupling between the metric $g_{ab}$ and the reference frame $\Re_{g}$.
In the following, to illustrate what I term the View from Nowhere, I examine an alternative type of reference frame. Specifically, I examine two distinct sets of GPS reference frames, which are identified as dynamical reference frames (DRFs) in Bamonti (2023), that is, reference frames dynamically coupled to the metric but without backreaction. For the present purposes, each GPS reference frame can be treated as a set of four scalar fields, corresponding to the proper time signals transmitted by four satellites. These signals, originating from a fixed initial point $O$, are transmitted to a target point $P$, effectively assigning four numerical values that ’coordinatise’ $P$. For a detailed account of the construction of a GPS reference frame, see Rovelli (2002a).
Building on the framework of Bamonti and Gomes (2024a), consider that the two sets of satellites define a red frame $\{\phi_{r}^{(I)}\}$ and a blue frame $\{\phi_{b}^{(I)}\}$, respectively. These frames represent two distinct ”physical parametrisations” over a shared spacetime region. Importantly, each reference frame is derived from a distinct physical system. In this context, the general-relativistic model can described by the tuple $\langle\mathcal{M},g_{ab},\phi_{r}^{(I)},\phi_{b}^{(I)}\rangle$, and is supplemented by initial data $(\Delta^{g},\Delta^{\phi_{r}}_{g},\Delta^{\phi_{b}}_{g})$ which specify the initial conditions for the metric and the two frames. As in equation (1), both $\phi_{r}^{(I)}$ and $\phi_{b}^{(I)}$ constitute local diffeomorphisms $U\subset\mathcal{M}\to\mathbb{R}^{4}$.
Using the frame-bundle formalism introduced earlier, the model space is $M:=(\mathrm{Lor}(\mathcal{M})\cup\Phi)$, where $\Phi=\{\phi_{r}^{(I)},\phi_{b}^{(I)}\}$ represents the space of GPS scalars defining the red and the blue frames. This model space is structured as a principal fibre bundle with the structure group $\mathrm{Diff}(\mathcal{M})$ and base manifold $[M]$. Here, $[M]$ comprises equivalence classes of metrics and reference frame, expressed as $\{[m]\in M\}=\{[g_{ab}]\in[\mathrm{Lor}(\mathcal{M})],[\phi_{r/b}^{(I)}]\in[\Phi]\}$.
If $g_{ab}$ satisfies the condition described in Equation (3) for some $F_{\phi_{r}}$ (resp. $F_{\phi_{b}}$), then any pair $(g_{ab},\phi_{r}^{(I)})$ (resp. $(g_{ab},\phi_{b}^{(I)})$), can be mapped into a unique reference frame representation of the metric. This mapping is achieved through the action of $f_{\phi_{r}}$ (resp. $f_{\phi_{b}}$), yielding $(g_{ab})_{F_{\phi_{r}}}:=f_{\phi_{r}}^{*}g_{ab}$, which is equivalent to the local gauge-invariant observable $[\phi_{r}^{-1}]^{*}g_{ab}:=g_{IJ}(\phi_{r})$. The same is for for $\phi_{b}$. Refer to Figure 2.
Figure 2: The space of models $M=\mathrm{Lor}(\mathcal{M})\cup\Phi$ with its gauge group $\mathrm{Diff}(\mathcal{M})$. Each point corresponds to a triple $(g_{ab},\phi_{r}^{(I)},\phi_{b}^{(I)})$. A reference frame (either $\phi_{r}$ or $\phi_{b}$) selects a unique representative $g_{IJ}(\phi_{r})$ (or $g_{IJ}(\phi_{b})$) for each fibre $\mathbb{F}_{m}$. This is achieved through the projection map $f_{\phi_{r}}$ (or $f_{\phi_{b}}$). The map projects a model within a fibre onto the intersection of the fibre with a chosen reference frame, represented by the ‘level surface’ $F_{\phi_{r}}$ (or $F_{\phi_{b}}$).
As per Equation (2), $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ are gauge-invariant, relational observables that represent distinct physical scenarios. Nonetheless, they are frequently interpreted as mere perspectives on a shared equivalence class $[g_{ab}]$ of isometric metrics. This equivalence class is regarded as the sole structure with ‘ontological significance’, which means that it is considered the fundamental structure underlying physical reality. [^22] To be blunt, this ontological picture, termed ‘Cartesian-Hegelian’ by Adlam (2024b), is not the same as that advocated by the moderate perspectivalism, which in contrast denies the claim that only the perspective-neutral picture, here represented by $[g_{ab}]$ is ontologically fundamental. According to moderate perspectivalism, “both perspectival and perspective-neutral facts exist and both are equally fundamental” [my emphasis] (ivi, p.9). However, my criticism of perspectivalism and the ontological fundamentality of $[g_{ab}]$ also extends to the moderate perspectivalism advocated by Adlam. In particular, I criticise the presence of the term ‘both’ on Adlam’s quote, in favour of the View From Everywhere, whereby equivalence classes $[g_{ab}]$ have no ontological status at all. Perspective-neutral facts exist in the same way that the number 5 exists. Note that the fact that moderate perspectivalism also recognises ontological fundamentality to the perspective-neutral structure supports the claim made in the introduction that moderate perspectivalism and the View From Nowhere are not at the antipodes. Both views support the ontological fundamentalism of equivalence classes. Remember the mantra: perspectives are always perspectives on something. As shown in Table 1, $[g_{ab}]$ is frame-free, spatiotemporally implicit, and gauge-invariant. This interpretation constitutes what I call:
The View from Nowhere: Frame representations are perspectives on an equivalence class of models implicitly defined without reference frames. This could also be termed a ”view from no-reference frame.” [^23] This interpretation differs from that of Adlam and Rovelli (2023), who associate a View from Nowhere with observer-independent facts, equating observers to reference frames. Here, following Wallace (2019), I assert that the View from Nowhere implies the existence of frame-free facts, distinguishing between observer-independence and observer-freedom.
Advocates of this view argue that physical reality fundamentally consists of abstract equivalence classes, existing independently of specific reference frame representations. These representations, which vary across spatiotemporal frames, are simply different ways of characterizing the same invariant structure. Consequently, transformations between reference frames are interpreted as notational changes rather than changes to the underlying reality.
Drawing from Kantian ontology (Kant, 1998), I differentiate between phenomenal reality, —measurable phenomena as they appear to observers— and noumenal reality, which represents the ultimate essence of reality. Within this framework, frame representations serve as phenomenal depictions of an implicit underlying reality, expressed through equivalence classes that cannot be explicitly characterised in spatiotemporal terms. In this terminology, the View from Nowhere presupposes the existence of a noumenal reality for which different spatiotemporal, phenomenological realities can be provided. [^24] This noumenal reality resonates strongly with Adlam (2024b)’s claim that perspective-neutral reality is not empirically accessible (so, not phenomenological in Kantian’s terms), but it is there. Explicitly, she states: “Although these [perspective-neutral] facts do not directly describe possible experiences […] they are physically real” (ivi, p.7).
While theoretically consistent, the View from Nowhere presents challenges for experimental practice. This view relies on an abstract, frame-free ontology that does not directly correspond to the relational nature of empirical data, which inherently depend on specific reference frames. As a result, translating theoretical constructs into measurable phenomena becomes a significant obstacle.
The relational character of empirical measurements has been widely acknowledged throughout the literature. Observations rely on the relationships established between the system under investigation and the reference frame within which measurements are conducted. For instance, Anderson emphasises that all measurements fundamentally involve comparisons between different physical systems (Anderson, 1967, p.128). Similarly, Rovelli (Rovelli, 1991, p.298) and Landau (Landau and Lifshitz, 1987, p.1) stress the indispensable role of reference frames in any measurement process. This reliance on reference frames underscores the difficulty of reconciling the abstract, frame-free ontology of the View from Nowhere with practical empirical methodologies. [^25] Additional support for the relational nature of empirical data comes from the ”Unobservability Thesis,” which posits that symmetry-related models of a system are empirically indistinguishable (Wallace, 2022c). Similarly, discussions on empirical (in)equivalence argue that explicit inclusion of observers leads to an ‘immanent’ conception of empirical distinctions, where models differ only if field configurations exhibit relational differences (Pooley and Read, 2021).
This operational stance aligns with Einstein’s original articulation of the point-coincidence argument, where he asserted that the physical content of a theory lies in the spacetime coincidences of material points (see Giovanelli (2021)). Specifically, Einstein highlighted that spacetime verifications invariably amount to determining such coincidences Einstein (1916) and that physical experiences are always assessments of point coincidences Einstein (1919). These arguments reinforce the relational nature of observations and challenge the View from Nowhere’s reliance on frame-free ontology.
Thus, any observable quantity, being both measurable and predictable, must be gauge-invariant and relational. By contrast, assigning ontological significance solely to abstract equivalence classes introduces significant challenges for observational verification, raising practical challenges for experimental validation. To address these issues, advocates of the View from Nowhere must go beyond asserting frame-free gauge-invariant content and provide practical guidance for experiment-based predictions—a daunting task, in my opinion.
The philosophical foundations of The View from Nowhere: Sophistication
Despite various criticisms, the View from Nowhere continues to enjoy a substantial following. Its conceptual basis is the so-called Sophistication approach to symmetries, which sharply contrasts with Eliminativism (or reductionism).
According to Eliminativism, a theory should be reformulated exclusively in terms of its symmetry-invariant quantities — typically equivalence classes — and thus eliminating all redundant, symmetry-related models.
In contrast, Sophistication adopts a structuralist stance by retaining the full set of symmetry-related models and treating them as isomorphic representations of a single, underlying invariant structure, which holds genuine ontological significance. Indeed, as Jacobs (2021) (drawing on Klein’s Erlangen Programme (Klein, 1893)) explains, Sophistication is best characterised as a symmetry-first (or external) approach. For further discussion of these positions, see, e.g., Dewar (2019) and Gomes (2023b).
A central debate in this context concerns the justification for treating symmetry-related models as physically equivalent. Two interrelated strategies have been proposed (Møller-Nielsen, 2017).
The interpretational approach insists that physical equivalence can only be granted once a clear and explicit (i.e., perspicuous) account of the invariant ontology is provided.
In contrast, the motivational approach is less demanding ontologically and adopts a more pragmatic stance. Rather than requiring an exhaustive intrinsic description of the shared structure, it demands only a compelling justification for treating the models as equivalent.
In this sense, Eliminativism aligns with the interpretational approach by seeking to eliminate all non-invariant features, whereas Sophistication typically resonates with the motivational approach by justifying equivalence without necessitating full intrinsic characterisation of the invariant structure.
However, proponents of Sophistication, notably Dewar, contend that Sophistication does not require an intrinsic characterisation of the space $[M]$ (the collection of equivalence classes) in order to secure ontological commitment to its elements. He maintains that the symmetries themselves—by revealing the invariant structure—suffice for ontological commitment in line with the interpretational approach. However, this position has not gone unchallenged. Critics such as Martens and Read (2020) argue that the sophistication strategy, when understood in this light, is ‘cheap’ because it fails to provide a perspicuous account of the shared invariant ontology underlying the various models.
In response to this critique, Gomes (2024a) defends Sophistication through the use of reference frames—termed representational conventions—which are formalised by a projection operator $f_{\sigma}:M\to M$ on the fibre bundle $M$. This procedure provides a perspicuous, yet choice-dependent, characterisation of $[M]$. The adoption of $f_{\sigma}$ in lieu of the section map $\sigma:[M]\to M$ supports the Sophistication claim that an intrinsic parametrisation of elements $[m]\in[M]$ is unnecessary. Furthermore, Gomes’ approach seeks to integrate aspects of both the interpretational and motivational approaches, shielding Sophistication from accusations of being ‘cheap’ and providing a perspicuous understanding of the ontology. That is, Gomes succeeds in demonstrating that answering why symmetry-related models represent the same physical reality, as required by the motivational approach, also gives us an insightful characterisation of ontology, as necessarily required by the interpretational approach.
Nonetheless, I argue that Gomes’ approach, while insightful, is ultimately insufficient. Even if one characterises each equivalence class by selecting a ’representative’ (a relational gauge-invariant observable $f_{\sigma}^{*}(\bullet)$), this does not justify the underlying ontological commitment to the equivalence classes themselves. In effect, designating a representative only offers one perspective on the broader structure, leaving the full ontology of the theory inadequately (or insufficiently perspicuously) addressed.
Loss of information
I have already expressed many of my qualms about the View from Nowhere, and in the next section I will offer an alternative on how to interpret the formalism of relational observables. Before doing so, however, I want to dwell on a further problem related to the View from Nowhere: namely the loss of physical information. [^26] This loss is exemplified by what Tong says about the choice of a gauge: The [gauge] redundancy allows us to make manifest the properties of quantum field theories, such as unitarity, locality, and Lorentz invariance, that we feel are vital for any fundamental theory of physics but which teeter on the verge of incompatibility. If we try to remove the redundancy by fixing some specific gauge, some of these properties will be brought into focus, while others will retreat into murk. By retaining the redundancy, we can flit between descriptions as is our want, keeping whichever property we most cherish in clear sight. (Tong, 2018, p.1) It can be said that, within this view, different choices of reference frames can be conceptualised as ‘windows of knowledge that provide partial views of a shared invariant structure’, available to an observer (see Adlam (2024a) on how to schematise a conscious observer in a diff-invariant theory such as GR). Accordingly, an observer can adjust her perspective to explore specific aspects of this invariant structure, effectively selecting one reference frame over another based on her investigative focus.
However, this approach implies a loss of the complete, ’absolute’ information contained within the ontologically fundamental structure, which is defined by the equivalence class.
This idea also resonates with Einstein (1917)’s assertion that “ […]a definite choice of the system of reference […] is contrary to the spirit of the relativity principle.” Similarly, (Adlam, 2024a, p.9) argues that “diffeomorphism invariance could finally be broken by the observer herself.” Geng (2024) also echoes this viewpoint.
Selecting a specific reference frame, such as through gauge-fixing, highlights certain properties of the physical system while concealing others. It is only by retaining redundancy and adopting a perspective from nowhere that the entirety of the physical landscape (the equivalence class) be fully represented. The View from Nowhere aspires to achieve this comprehensive perspective, though it does so at the expense of direct empirical applicability.
This critique of the View from Nowhere highlights significant conceptual and practical challenges, though further exploration may reveal additional nuances. In the following section, I will explore how to provide a perspicuous characterisation of the theory’s invariant ontology without resorting to equivalence classes of symmetry-related models, as the sophisticated approach advocates. I shall refer to this alternative perspective as the View from Everywhere.
4 The View from Everywhere: Frame representations are all that exist
Returning to the example of the two GPS reference frames introduced earlier, an alternative interpretation of the two local observables $g_{IJ}(\phi_{b})$ and $g_{IJ}(\phi_{r})$ can be proposed. Rather than treating frame representations as perspectives on an abstract equivalence class, this alternative asserts that each member of the collection of observables $\{g_{IJ}(\phi_{b}),g_{IJ}(\phi_{r}),\dots\}$ (where the ellipsis indicates additional reference frames and related observables) constitutes all that fundamentally exists.
This perspective embodies what I term:
The View from Everywhere: each frame representation is all that ultimately exists. It may also be conceptualised as a “view from every reference frame”, emphasising the primacy of local, relational observables.
In Bamonti and Gomes (2024b), a map m called external diffeomorphism is introduced to relate the two frames $\{\phi_{r}^{(I)}\}$ and $\{\phi_{b}^{(I)}\}$, which function analogously to a coordinate transformation. Unlike an ‘ordinary diffeomorphism’, an external diffeomorphism acts directly on the already constructed local, gauge-invariant observables, changing frames and getting us to a different and new observable.
By redefining $\phi_{r}^{(I)}:=X^{I}_{r}$ and $\phi_{b}^{(I)}:=X^{I^{\prime}}_{b}$, the gauge-invariant observables $g_{IJ}(\phi_{r}):=g_{IJ}(X^{I}_{r})$ and $g_{IJ}(\phi_{b}):=g_{I^{\prime}J^{\prime}}(X^{I^{\prime}}_{b})$ are connected via the map m, which operates as follows:
$\text{\bf{m}}:g_{IJ}(X^{I}_{r})\rightarrow g_{I^{\prime}J^{\prime}}(X^{I^{\prime}}_{b})=\cfrac{\partial X^{I}_{r}}{\partial X^{I^{\prime}}_{b}}\cfrac{\partial X^{J}_{r}}{\partial X^{J^{\prime}}_{b}}g_{IJ}(X^{I}_{r}).$ (5)
Clearly, this represents a passive diffeomorphism transformation. [^27] Due to the one-to-one correspondence between active and passive diffeomorphisms, from the active perspective, the relationship between $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ can be interpreted as the external diffeomorphism $\textbf{d}:=\phi_{r}^{-1}\circ\textbf{m}\circ\phi_{r}$, or equivalently $\textbf{d}:=\phi_{b}\circ\textbf{m}^{-1}\circ\phi_{b}^{-1}$. This shows that in GR, local gauge-invariant observables are covariant under frame transformations [^28] Pitts (2022) refers to covariance of observables relative to coordinate systems. Here, I propose to extend this notion to encompass reference frames., so the introduction of a gauge choice (in the form of a choice of reference frame) does not spoil either the gauge invariance or the covariance of the theory.
As I will further stress below, it is important to note that there can indeed be frame-independent, explicit observables, however these quantities are inherently non-local in nature (see Table 1 and my notion of Relality below).
While the map m provides a ’shared vocabulary’ for translating between two distinct reference frame representations, it is important to emphasise that these represent separate and fully-fledged physical situations. [^29] In (Belot, 2017, p.954)’s terminology, m is a “physical symmetry”—an isomorphism linking solutions that represent distinct “possibilia”—as opposed to a “gauge symmetry,” which relates solutions that cannot be taken to represent distinct physical states. Consequently, they are not two perspectives of a shared, total physical state as proposed in the View from Nowhere. Unlike the View from Nowhere, the View from Everywhere does not require the existence of such a shared state. Instead, each gauge-invariant observable represents an independent and self-contained physical reality. Under this framework, we do not require any invariant, frame-free structure in our ontology. Instead, the focus shifts to a theory of frame-dependent yet gauge-invariant objects.
Importantly, the absence of a shared, total reality in this framework—replaced by a collection of local, frame-dependent realities—does not imply a lack of coherence. Fragmentation does not imply incoherence. The different physical situations represented by the gauge-invariant observables $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ remain interconnected through external diffeomorphisms m.
The map m also protects us from potential accusations of solipsism (see (Adlam and Rovelli, 2023, sec.3) and the references therein for a discussion in the context of relational quantum mechanics). Each observer—understood as being associated with a reference frame—is not isolated in their representation of reality but can communicate with all other observers. Through m, observers can translate their frame-specific representations into those of others. Importantly, this does not mean that each observer accesses merely a fragment, a perspective of a larger, overarching whole. Instead, every perspective constitutes a complete and self-consistent depiction of reality.
Furthermore, it is worth noting that the map m and the possibility of inter-frame communication is a frame-independent fact. This does not contradict the principles of the View From Everywhere, which do not exclude the possibility of frame-independent facts but only the possibility of frame-free facts. This last observation is fundamental to the introduction of the notion of Relality, which is based on the View From Everywhere. [^30] This flexibility of the View From Everywhere also protects it from the so-called ‘iteration of relativisation’ attacks reported in the literature, e.g. in Riedel (2024), from which similar relational approaches are said to suffer.
Relality
The existence of the external diffeomorphism m enables advocates of the ”View from Everywhere” to define an equivalence class $[g_{IJ}]:=\{g_{IJ}(\phi_{r}),g_{IJ}(\phi_{b}),\dots\}$, which differs fundamentally from the equivalence class $[g_{ab}]:=\{g_{ab},(d^{*}g)_{ab},\dots\}$ used in the View from Nowhere. Unlike the latter, the equivalence class $[g_{IJ}]$ within the View from Everywhere supervenes on the ensemble of all the frame-dependent quantities. Each of these quantities form the a bona-fide physical reality, rather than existing as perspectives of a ‘more complete’, frame-free structure.
Coherently, $[g_{IJ}]$ is a frame-independent object, as it is expressed across all possible reference frames. So differently from $[g_{ab}]$, it is not a frame-free object. In formal terms: the View from Everywhere retains a one-to-one correspondence between the equivalence class $[g_{IJ}]$ and the orbit $\mathcal{O}_{g}:=\{g_{IJ}(\phi_{r}),g_{IJ}(\phi_{b}),\cdots\}$. However, unlike the View from Nowhere, $[g_{IJ}]$ in the View from Everywhere does not exist independently of the frame-specific representations within the space of models. Essentially, $[g_{IJ}]$ and $\mathcal{O}_{g}$ coincide. See Figure 3.
Figure 3: The set $[M]$ of equivalence classes is part of the space of models itself. Moreover, it is the set of relational observables that can be understood as a fibre, generated by the group $Diff(\mathbb{R}^{4})$, whose group elements are external diffeomorphisms m.
If for the View from Nowhere the equivalence class $[g_{ab}]$ represented a shared ontology, for the View from Everywhere, the equivalence class $[g_{IJ}]$ only represents merely the presence of a shared vocabulary. This distinction underscores the divergence between the two views: the View from Everywhere rejects the existence of a frame-free structure, focusing instead on the relational and frame-dependent nature of observables. Consequently, the term ’perspective’ becomes problematic in this context, as it implies an underlying, autonomous reality that can be described from multiple viewpoints—an assumption the View from Everywhere does not adopt.
What conception of reality emerges from this framework? Within the View from Everywhere, each of the frame-dependent physical situations constitutes a fully-fledged reality. The collection of such local realities give rise to what I call Relality, which however is not an ontologically independent structure of individual frame-dependent representations: it is only a name that indicates the possibility of inter-translatability of frame-dependent and local fundamental realities.
The concept of Relality is intriguing and deserves to be analysed further. Importantly, this concept helps me to emphasise that reality is fundamentally local and relational.
Drawing on the Kantian perspective used to interpret the View from Nowhere, I argue that the View from Everywhere fundamentally rejects the notion of a shared, implicit noumenal reality underlying its phenomenological manifestations. [^31] This reference to phenomena is in no way related to French (2023)’s phenomenological approach. Here, I make no reference to the cognitive role or knowledge of conscious observers. An observer is understood as any (even non-conscious) physical system and ‘what appears to him’ is understood simply as what he enters into relation with through some physical interaction. Thus, the phenomenon is understood as the relational empirical content as encoded within the partial/complete observables framework. According to the View from Everywhere, there is no ’hidden’ structure beyond these frame-specific representations; each gauge-invariant quantity constitutes the entirety of what fundamentally exists. Accordingly, the two observables $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ do not describe two distinct phenomenal perspectives of a shared, ontologically independent noumenal reality represented by $[g_{ab}]$. Instead, each constitutes a local reality and there is nothing else. This is because the concept of Relality, formalised by $[g_{IJ}]$, does not represent a shared ontology: it is not an ontological structure ‘beyond’ and in addition to realities $\{g_{IJ}(\phi_{r}),g_{IJ}(\phi_{b}),\dots\}$. The definition of $[g_{IJ}]$ only ensures the inter-translatability of each local reality. [^32] One should not confuse this definition of Relality with the so-called first-person-plural view typical of epistemic perspectivalism. For in that view, the common vision of the community is broader and more complete than the individual perspectives on that vision. By contrast, according to my concept of Relality, the whole is not ‘something more’ than the individual parts. Relality only encompasses the interconnectedness of complete realities, but there is no collective, total vision to be discovered or achieved: neither epistemically nor ontologically. In this regard, the notion of Relality should also not be understood as confirming or disconfirming the legitimate presence of frame-independent, explicit and non-local observables. It only concerns local, relational observables. This clarification is important: it also prevents the possible attack on the frame-independent m-map according to which the exchange of information via connections between frames cannot be defined with respect to a frame-independent collective structure ($[g_{IJ}]$ precisely), which presupposes the existence of such connections in order to be defined (see e.g. Adlam, 2024b, §4.3). The map m is frame-independent in the sense that it is defined relative to any observer: there is no need for a collective structure relative to which it is defined. All of this supports the abandonment of the concept of ‘perspective’.
I argue that the interpretive framework of the View from Everywhere challenges the core assumptions of the Sophistication approach. Specifically, I will demonstrate that the View from Everywhere offers greater ontological parsimony—minimising unnecessary metaphysical commitments—and provides a clearer, more perspicuous characterisation of the ontology underlying relational observables. Consequently, I contend that Sophistication is neither the ultimate nor the most effective framework for interpreting relational observables.
Beyond Sophistication
As already stressed, the Sophistication approach posits that the fundamental description of physical reality should rely on equivalence classes. The View from Everywhere proves that this assumption is unnecessary. Gauge-invariant relational observables, such as $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$, represent distinct, fully-fledged physical scenarios, leaving no basis to hypothesise the existence of an additional ’hidden’ structure beyond them. This interpretation aligns with a principle of ”ontological parsimony”, which I sustain. The View from Everywhere, being ontologically more conservative, posits no ‘entities’ beyond the realities expressible within reference frames.
As discussed previously, adopting Sophistication while supporting a motivational approach presents significant challenges, particularly when attempting to reconcile it with the demand for a perspicuous characterisation of ontology, which is more closely tied to the interpretational approach. Among the proposals, Gomes’ motivational approach shows promise but ultimately falls short within the Sophistication framework. However, Gomes’ motivational approach aligns naturally with the View from Everywhere. By eliminating the need for a shared ontology based on equivalence classes, the very construction of the gauge-invariant observables constituting the fundamental ontology inherently provides a perspicuous characterisation of the ontology.
This claim raises a delicate question: How can Gomes’ approach, which employs projection maps onto fibres of isomorphic, gauge-variant models, function within the framework of the View from Everywhere? After all, I have argued that the only relevant fibre in this framework consists of gauge-invariant observables connected by external diffeomorphisms (see Figure 3). At first glance, it appears that the View from Everywhere, if not based on Sophistication, instead relies on Eliminativism, rendering Gomes’ proposal inapplicable. However, this conclusion is incorrect. The View from Everywhere accommodates the formal construction of frame representations within the fibre bundle formalism, demonstrating the continued relevance of Gomes’ insights. For example, a frame representation, such as $g_{IJ}(\phi_{r})$, can still be derived as a projection of a fibre onto a section. Even if within the View from Everywhere, the ontology is characterised by quantities that are independent of the structures distinguishing various isomorphic, gauge-variant models, the construction of these quantities necessitates the use of non-invariant objects. This fact lies at the heart of the relational observables strategy, where gauge-variant quantities serve as handles through which other gauge-variant quantities are coupled, to form gauge-invariant quantities (see Rovelli (2014)).
Therefore, supporting View from Nowhere does not imply the removal of redundant gauge degrees of freedom. Rather, it only involves attributing ontological importance to the gauge-invariant observables.
No loss of information
Previously, I outlined one advantage of the View from Everywhere: its ability to provide a perspicuous characterisation of the ontology of the theory, in line with a motivational approach. In this concluding paragraph, I emphasise an additional advantage: it resolves the issue of physical information loss and the problem of diffeomorphic symmetry breaking, both of which stem from the choice of a particular reference frame (or, equivalently, gauge fixing).
First, the breaking of the diffeomorphic freedom of theory is precluded by the existence of the external diffeomorphism m. This map relates different reference frame choices333333m serves as the analogue of the transition map between sections of a principal bundle, given by $\mathrm{\bf{t}}_{\sigma\sigma^{\prime}}(\varphi)=f_{\sigma^{\prime}}(\varphi)^{-1}f_{\sigma}(\varphi)$ (see Gomes (2024a))., ensuring GR to remain m-covariant (see also Bamonti and Gomes (2024a)).
Second, the notion of Relality implies that all the available information resides within each frame. Consequently, there is no ‘total’ information that could be lost when a specific reference frame is selected. This may also alleviate the concerns raised in Wallace (2024) that the choice of certain reference frames (some gauge-fixings) does not preclude the physical features of the system from varying under a gauge transformation, leading to a loss of genuine information and casting doubt on whether they can really be classified as gauge-fixing at all (see ‘The Incompleteness Reason’, ivi, p.13). [^34] To be precise, Wallace studies the so-called ‘unitary gauge’ in electrodynamics. I argue that this choice of gauge corresponds to the use of uncoupled reference frames. In other words, the physical features that change are the features of the physical system considered isolated from the environment (Wallace, 2022a, b). This isolation procedure coincides with the dynamic uncoupling of the reference frame. The “loss of genuine information” referred to by Wallace, therefore, is not really genuine because it is a consequence of this approximate procedure. Within the View from Nowhere, on the other hand, there is the loss of bona-fide genuine information as a result of the use of coupled reference frames, which identify distinct gauge-invariant observables. The View from Nowhere alleviates this more serious problem.
Before concluding, I wish to highlight another advantage of the ”View from Everywhere” concerning the choice of reference frame. The choice is entirely arbitrary, raising the question of what guides such a choice. In the context of the View from Nowhere, reference frame selection is typically driven by pragmatic considerations, aimed at providing a convenient description of an objective, overarching physical reality. These motivations are largely ’conventional,’ reflecting a preference for ease of description rather than any deeper ontological commitment. In fact, the choice of reference frame, by definition, merely offers a conventional, perspectival description. [^35] A contrasting argument is discussed in (Gomes and Butterfield, 2024, sec.3). In the context of electromagnetism, the authors propose that “a choice of gauge need not be a matter of calculational convenience for some specific problem or class of problems, but can be related to a physically natural and general splitting of the electric field.” However, the authors also note that a physically defined gauge choice (e.g., the Coulomb gauge, which splits the electric field into radiative and Coulombic parts, with the latter determined by the instantaneous charge distribution) is still “non-mandatory.” It corresponds to a particular choice of electric field decomposition. Thus, ultimately, it remains, in my view, a “convention.” In contrast, in the context of the View from Everywhere, a choice of reference frame constitutes an empirical choice, rather than a conventional one. This is because we cannot conventionally choose one or another reference frame that represents the ‘same’, ‘true’ physical situation in the most helpful way for the purpose at hand. A different choice is in all respects a different reality.
Finally, the View from Everywhere also alleviates the problem whereby the validity of a choice of reference frame cannot be empirically tested. In fact, every measurement is always made within a reference frame, i.e. to make measurements, a reference frame must necessarily be set. Thus, we have no ‘meta-empirical ground’ (that is, a frame-free empirical ground) to compare the various possible reference frames. Within the View from Nowhere, this is a problem, since each choice of reference frame is comparable to a perspective and it makes sense to ask which description is ‘the best’, from a pragmatic or fundamental point of view. The advocate of the View from Everywhere, on the other hand, might claim that there is no need to test the validity of the choice of reference frame, since any choice of reference frame is inherently ’the best’ or ’the most fundamental’, as the only available and existing reality.
5 Conclusion
This paper has employed the fibre bundle framework to examine the role of reference frames and relational observables in GR, highlighting the philosophical and practical consequences of two contrasting paradigms: the View from Nowhere and the View from Everywhere. Below, I summarise the key findings of the paper and their broader implications.
In Section 2, I used the fibre-bundle formalism to discuss reference frames and relational observables in GR. This approach defines the space of models as a principal bundle and introduces gauge orbits to describe symmetries within the theory. Reference frames, acting as section maps, facilitate the construction of gauge-invariant observables. This formalism elucidates the relationship between the choice of a reference frame and the choice of a gauge-fixing.
In Section 3, I introduced the View from Nowhere, which conceptualises reference frame representations as perspectives on an underlying invariant equivalence class of isomorphic models. This perspective, informed by the Sophistication approach, posits that the fundamental ontology of GR resides in gauge-invariant equivalence classes, independent of specific frame representations. While theoretically robust, this perspective encounters significant challenges in aligning with empirical practice, as it assumes the existence of an abstract, frame-free reality that is not directly observable.
In Section 4, I introduced the View from Everywhere. Contrary to the View from Nowhere, it rejects the presumption of an underlying equivalence class to be given ontological relevance. Instead it asserts that frame-dependent, local relational observables constitute all that fundamentally exists. To capture this viewpoint, the concepts of Relality and of external diffeomorphism m were introduced, emphasising the fragmented yet coherent nature of frame-dependent realities. By adhering to the principle of ontological parsimony, the View from Everywhere avoids unnecessary metaphysical commitments to frame-free structures and naturally provides a perspicuous characterisation of the ontology, with implications for experimental design and theoretical consistency.
The philosophical implications of adopting the View from Everywhere warrant further examination. This includes analyising concepts like perspectivalism and objectivity within GR. A deeper analysis of the concept of ’Relality’ may yield valuable insights in this regard. Additionally, investigating the role of external diffeomorphisms within the context of quantum reference frames offers a promising direction for future research.
In conclusion, this paper has examined the complex interplay between reference frames, gauge symmetries, and ontology in GR. By contrasting the View from Nowhere and the View from Everywhere, I have sought to deepen our understanding of the interplay between theory, empirical practice, and philosophical interpretation. This analysis underscores the significance of reference frames in shaping our understanding of physical reality and highlights the need for further investigation into their foundational role in physics.
Acknowledgement
I thank Henrique Gomes for useful discussions. I thank S. for inspiring me to conceive the View from Everywhere.